What is the area enclosed by the graph of

y = x2 + x – 2

 below the x-axis?

 


1
Expert's answer
2021-06-07T14:59:06-0400

we have equation

"y = x^2 + x \u2013 2"

we have to get the area below the x-axis ,see the graph below to get better idea


this part of parabola is required area

so we have to solve the integral "\int_{-1}^{2}x^2-x+2 dx
\\"

"=[\frac{x^3}{3}-\frac{x^2}{2}+2x]_{-1}^{2}=\frac{-9}{2}=-4.5"

taking mod the answer is 4.5 unit"^2"